OpenStudy (anonymous):

Please help! What is the slope of the line which passes through (–2, 0) and (0, 4)? –2 0 2 Undefined

OpenStudy (ghazi):

slope = (4-0)/(0-(-2)) =4/2=2

OpenStudy (ghazi):

\[m= \frac{ (Y _{2}-Y _{1}) }{ (X _{2}-X _{1}) }\] m-- slope